Lucinda has a 5.6‑mH inductor. What size capacitor ? should she choose to make an oscillator with a natural frequency of 143 kHz?
Answer #1

Oscillator with natural frequency is given by:

When

Capacitive reactance = Inductive reactance

XL = XC

w*L = 1/(w*C)

2*pi*f0*L = 1/(2*pi*f0*C)

C = 1/(4*pi^2*f0^2*L)

So, Using given values:

C = 1/(4*pi^2*(143*10^3)^2*5.6*10^-3)

C = 2.2*10^-10 F = Capacitance of capacitor

Please Upvote.